OpenStudy (hayleymeyer):

The volume of a cylindrical can is 15.7 in.3. The diameter is 2 in. Find the height. Use 644-12-03-00-00_files/i0110000.jpg = 3.14. a. 1 in. c. 5 in. b. 62.8 in. d. 3.14 in.

OpenStudy (hayleymeyer):

use pie=3.14

OpenStudy (hayleymeyer):

@Ap3xPred

OpenStudy (anonymous):

The volume of a cylinder is the height times the area of the base. The area of the base is pi times the radius squared. The radius is half the diameter. Therefore the volume can be described as: v = h*pi*(d/2)^2 Plugging in the given values gives us: 15.7 = h*3.14*(2/2)^2 Now simplify and solve for h: 15.7/3.14 = h*1^2 = h h = 5 inches

OpenStudy (hayleymeyer):

thanks!

OpenStudy (anonymous):

you're welcome